Which of the saturated solutions below would have the highest [OH-1]? Which would have the lowest pH?

M(OH)2
Ksp= 4.25*10^-6
Ksp= 7.39*10^-4
Ksp= 2.64*10^-3
Ksp= 8.52*10^-8
Answer #1
highest [OH]- = for Ksp= 2.64*10^-3

lowest pH = Ksp= 8.52*10^-8
Answer #2
a)highest [OH-1] is Ksp= 2.64*10^-3
b)lowest pH is Ksp= 2.64*10^-3
explanation:
Ksp = [OH-1]^2*K
and pOH is antilog of [OH-1]
pH is 14-pOH